3/4x – 5= 2/3x: 3/4x – 5= 2/3x
Answer by jim_thompson5910(9166)


(3/4)x-5=(2/3)x Start with the given equation



(12)((3/cross(4))x-5)=(12)((2/cross(3))x) Multiply both sides by the LCM of 12. This will eliminate the fractions (note: if you need help with finding the LCM, check out this solver)



9x-60=8x Distribute and multiply the LCM to each side



9x=8x+60Add 60 to both sides


9x-8x=+60 Subtract 8x from both sides


x=60 Combine like terms on the left side

--------------------------------------------------------------
Answer:
So our answer is x=60